# In-Context Learning Strategies Emerge Rationally

### v1

@@ 4.1 Novel Predictions
  • Sub-linear growth from generalizing to memorizing predictor. Given the empirical match between the Bayesian model’s posterior probabilities and the relative distance computed using the Transformer (see Fig. 6), we examine whether relative distance takes the functional form predicted by the model: sublinear growth with respect to $N$, and sigmoidal growth with respect to $N^{1-\alpha}$ (holding $D$ constant). This arises from Eq. 4, 5, wherein evidence accumulates sub-linearly with the number of samples $N$. As can be seen in Fig. 6(a), the relative distance of Transformer predictions between the memorizing and generalizing predictors clearly exhibits a sub-linear trend with $N$ and sigmoidal growth with respect to $N^{1-\alpha}$, in accordance with the functional form predicted by Eq. 4. Note also that the bottom left panel of Fig. 6(a) clearly shows that even at high task diversity values, relative distance slowly increases towards the memorizing predictor in a sigmoidal manner. This stands in contrast to Raventós et al. [25]’s claim that at high $D$ the Transformer only becomes closer to the generalizing predictor with increasing $N$ (see App. J for further discussion).
@@
  • Rapid behavior change near crossover. At the point where the two hypotheses have equal log-posterior odds, our model predicts there will be a crossover in which predictor dominates the posterior and explains model behavior. Given our sigmoidal functional form for $\eta(N,D)$, we can expect this change to be rapid—small variations in experimental conditions (e.g., task diversity) will yield large changes in model behavior. To confirm this, we plot second-derivative of relative distance in Fig. 6(b), finding its magnitude is indeed greatest at the boundary $\eta(N,D)=0$.
@@
  • Scaling of time to Transience with Task-Diversity. For a given value of $D$, we can predict the critical training time at which transience, i.e., the crossover from generalizing to memorizing predictor, occurs by solving for $N^{*}$ in $\eta(N^{*},D)=0$ (i.e., relative distance $d_{\text{rel}}=0.5$). Specifically, we have: $N^{*}(D)=\left[\frac{\Delta K(D)^{\beta}}{\gamma\Delta L(D)}\right]^{\frac{1}{1-\alpha}}$. We thus plot the predicted time to transience as a function of $D$ in Fig. 6(c). As can be seen, beyond the two hypotheses threshold (which determines the minimum $N$ value in which our model holds, see App. E), our predictions hold well with the empirically observed data. These results indicate super-linear growth of time to transience with task diversity. Importantly, we note that if this prediction derived from our model holds beyond our tested settings, it suggests that if the denominator in the expression becomes very small, e.g., when there are enough tasks in the mixture such that the memorizing and generalizing predictors produce similar outputs, the time to transience can approach infinity. In such a condition, generalization will persist regardless of how long a model is trained. Interestingly, we find that using a learning rate annealing schedule enhances the Transformer’s adherence to Bayes-optimal trajectories from generalization to memorization, and so we use it in this experiment. However, the effect of annealing appears to decay throughout training, yielding a slower rate of advancing toward memorization, which can be seen by the last two observed transience points veering from our model’s predictions (see App. K).
@@ 4.2 The Loss-Complexity Tradeoff
To further exemplify the intuition above, we can consider variations on the training pipeline that can be expected to affect a model’s ability to implement more complex predictors. Using the number of parameters as a control proxy to this end, we find that scaling MLP width raises the transition boundary between memorization and generalization—i.e., memorization is preferred more. Our model closely captures this effect: we see a decreasing complexity penalty $\beta$ with larger MLP width, pointing toward the possibility that larger models penalize complexity less, yielding more memorization. Interestingly, we find that $\beta$ decays exponentially with MLP width. Finally, we also vary context length and task dimensionality, and find that, once again, our model captures the effects of these interventions (see App. H.4).
@@ 5 Discussion
Here, we discuss the relation between our work and other Bayesian or normative accounts of ICL. We provide an extended review of related work in App. B. Several prior works framed ICL as Bayesian at inference-time, meaning models implement a single Bayesian predictor in context [35, 72, 50, 73, 25, 61]. The focus on a single predictor in these works often led them to view ICL as a single strategy, with some focusing solely on a memorizing solution [35], while others refer mainly to a generalizing solution as ICL [25]. While there are cases in which in-context behavior is well approximated by a single predictor (e.g., the memorizing solution for low $N,D$), our results robustly show that several predictors are required to fully capture model behavior, and that the extent to which a particular predictor explains model behavior varies across training. Thus, a posterior-weighted average over different predictors, which considers a bias towards different predictors coming from training, rather than only inference-time, is required to fully capture model behavior across conditions. In contrast with studies focusing on inference-time, two recent works, Carroll et al. [30] and Elmoznino et al. [71], offer Bayesian or normative views of pretraining to perform ICL. Importantly, while these papers take different theoretical perspectives from ours, they arrive at a similar conclusion as us: the existence of a tradeoff between loss and solution complexity in pretraining. Elmoznino et al. [71] offer a normative theoretical analysis of training to perform ICL via next-token prediction loss, showing it yields an Occam’s razor objective which minimizes loss as well as the learned solution’s complexity. Carroll et al. [30] study task diversity effects and transient generalization in the linear regression setting of Raventós et al. [25]. Their Bayesian account of pretraining, which is rooted in theory of singular models [74] and makes different assumptions from ours, interestingly yields a relatively similar functional form for the posterior-odds. However, their measure of complexity is architecture-dependent  [75], thus they only provide a qualitative analysis as they cannot directly estimate the complexity of Bayesian predictors. In contrast, our hierarchical Bayesian framework provides a quantitative, predictive account of pretraining phenomena, in addition to capturing inference-time behavior as a posterior-weighted average of solutions, which is not addressed by Carroll et al. [30]. Despite that, we view Carroll et al. [30] and Elmoznino et al. [71]’s works as valuable contributions, in particular with regards to offering potential explanations for the source of the simplicity bias seen in Transformers.
@@
While we rely on a specific theoretical abstraction to arrive at our results, i.e., a hierarchical Bayesian model, we believe the predictive power of this abstraction corroborates its faithfulness. However, one limitation of our analysis is use of settings where model behavior is largely explained by only two predictors. It can be interesting to analyze settings where more predictors are feasible, e.g., the four predictors identified by Park et al. [26] in their task of learning a finite mixture of Markov chains. We believe this extension would be straightforward: we currently only characterize the dynamics of learning a memorizing vs. a generalizing predictor, but additionally modeling the statistics used to produce outputs (bigram vs. unigram) should accommodate the solutions identified by Park et al. [26]. Additionally, our current modeling analysis focuses on capturing Transformer behavior for sequences seen during training (ID), and extending it to OOD evaluation is an important step. Finally, a crucial limitation of our analysis comes from the simple relation we assume between algorithmic complexity and complexity of implementation by a Transformer—while the assumption of simplicity bias is well-backed by theoretical and empirical claims [45, 44, 43], we believe our assumed relation could be improved by building on recent advances defining measures of how many effective parameters are used by a model to implement a solution [75, 76].
@@
We see two several takeaways from our work for understanding of deep learning.
@@ Acknowledgments
We thank the Computation and Cognition Lab, in particular Ben Prystawski and Michael Li; Jay McClelland, Satchel Grant, Jerome Han and the PDP group; the Physics of Intelligence group at Harvard, especially Eric Bigelow and Sonia Murthy; the CRISP lab at Harvard; Jesse Hoogland and Matthew Farrugia-Roberts; Surya Ganguli and the Neural Dynamics and Computation Lab; and Navin Goyal and the theory group at Microsoft Research India for useful discussions.
@@ Appendix A Glossary of Useful Terms
Generalizing Predictor ($G$). A predictor defined by a posterior-weighted average with a continuous prior over the true data-generating distribution $\mathcal{T}_{\text{true}}$. Such a predictor does not depend on the tasks seen during training, and can hence generalize well to novel, unseen tasks. This predictor maps onto ‘task-learning’ or ‘inference’ notions of ICL [26]. See also App. G for the precise mathematical form of generalizing predictors for settings studied in this work. Memorizing Predictor ($M$). A predictor defined by a posterior-weighted average with a discrete prior defined over the distribution of tasks seen by the model during training, $\mathcal{T}_{train}$. Such a predictor depends on the tasks seen during training, and hence generalizes primarily to those tasks. This predictor maps onto ‘task-retrieval’ notions, though it also maps onto ‘in-weights learning’ in the case of the classification task introduced by Chan et al. [20]. See also App. G for the precise mathematical form of memorizing predictors for settings studied in this work. Task diversity Threshold ($D^{*}$). For a given number of training steps $N$, task diversity threshold is defined as the critical number of tasks that must be involved in a training mixture such that the model behavior transitions from resembling a memorizing predictor to a generalizing predictor. Transient Generalization ($N^{*}$). For a given task diversity $D$, when a model is trained for sufficiently long, its behavior transitions from more closely resembling a generalizing predictor to more closely resembling a memorizing one. We define this point as the point in which the relative distance between memorizing and generalizing crosses $0.5$. This phenomenon is called transience, transient generalization, or the transient nature of in-context learning (see App. B for a longer discussion). The critical time to reach this point is denoted $N^{*}$ in the main paper. Relative Distance ($r$). A measure we use to characterize trade-off between predictors in settings where there are primarily two predictors (though one can easily generalize this measure to a setting with more predictors). Specifically, given a model $h$, two predictors $Q_{1},Q_{2}$, and a distance function $d$, we define the term $r:=\frac{d(h,Q_{1})-d(h,Q_{2})}{d(Q_{1},Q_{2})}$ and then relative distance as $d_{\text{rel}}=\nicefrac{{(r+1)}}{{2}}$. The latter operation rescales $r$ to a scale of 0 (if $h=Q_{1}$) to 1 (if $h=Q_{2}$), essentially assessing where the model $h$ lives when a line is drawn with endpoints ranging from $Q_{1}$ to $Q_{2}$. Posterior Odds, Prior Odds, Bayes Factor. Consider a set of observations $X$ and two hypotheses $H_{1}$ and $H_{2}$ that are being assessed as candidates to explain the data. Prior odds are defined as the ratio $\frac{P(H_{1})}{P(H_{2})}$, i.e., if no observations are seen yet, which hypothesis is apriori preferred. Bayes factor is defined as the ratio of likelihoods $\frac{P(X|H_{1})}{P(X|H_{2})}$, i.e., once the observations are received, we compare how likely individual hypotheses deem these observations. Finally, posterior odds are defined as the ratio $\frac{P(H_{1}|X)}{P(H_{2}|X)}=\frac{P(H_{1})}{P(H_{2})}\times\frac{P(X|H_{1})}{P(X|H_{2})}=\text{Prior odds}\times\text{Bayes factor}$, i.e., how do observations affect the prior to help assess which hypothesis is more likely.
@@ B.1 Prior Work Studying Task-Diversity Effects and Transience
We first discuss prior work studying phenomenology of ICL that forms the core of our paper: task diversity effects and transience. These works often focus on providing mechanistic accounts that are more bottom-up in nature, i.e., suitable for studying specific settings. For example, the induction head is a core mechanism employed by a model to perform the in-context classification task by Singh et al. [24], Reddy [28]. When transience occurs, the role of induction head is diminished, since it is not vital for implementing a memorizing predictor. However, as we show, transience occurs across a spectrum of settings, including ones where induction heads are never learned during training (e.g., Balls and Urns, which we analyze using a one-layer Transformer)—prior mechanistic accounts would not help justify transience in such settings. Our work thus takes a top-down account, i.e., we offer insights based on a computational model of ICL developed by capturing its phenomenology. This helps identify relevant control variables controlling a dynamic we argue lies at the heart of transience and task diversity effects: tradeoff between loss and complexity of a solution, which manifests itself into transitions between generalizing and memorizing predictors. Reconciling our computational model with mechanistic approaches undertaken in prior work can be an exciting avenue for progress.
  • Task diversity Threshold. A phenomenon first popularized by Raventós et al. [25] in an in-context linear regression task, albeit originally demonstrated by Kirsch et al. [60] in a prior work on in-context classification of permuted MNIST images and recently expanded to a Markov modeling setting by Park et al. [26], to another classification setting by Nguyen and Reddy [29], as well as to a modular arithmetic setting by He et al. [59]. Specifically, Raventós et al. [25] empirically show that if one trains Transformers on a mixture of linear regression tasks, there is a critical number of tasks below which the model behavior is well-characterized by a Bayesian predictor over the seen tasks (what we term the memorizing predictor), while above it the behavior is well-characterized by the standard solution of ridge regression. These results were expanded in a recent theoretical work by Lu et al. [78], who study the asymptotic dynamics of a linear attention Transformer and show the change in solution used by the model as a function of task diversity is a second-order phase transition. In contrast to their work, our empirical analysis covers a broad range of settings that involve sequence modeling, regression, and classification, while the analytical parts of our paper provide a predictive framework that identifies the relevant control variables and explains how they affect the behavior of standard, nonlinear Transformers across all studied settings.
@@
B.2 Broader Work on Understanding ICL and Meta-Learning

### v2

@@ 4.1 Novel Predictions
  • Sub-linear growth from generalizing to memorizing predictor. Given the empirical match between the Bayesian model’s posterior probabilities and the relative distance computed using the Transformer (see Fig. 6), we examine whether relative distance takes the functional form predicted by the model: sublinear growth with respect to $N$, and sigmoidal growth with respect to $N^{1-\alpha}$ (holding $D$ constant). This arises from Eq. 4, 5, where evidence accumulates sub-linearly with the number of samples $N$. As can be seen in Fig. 6(a), the relative distance of Transformer predictions between the memorizing and generalizing predictors clearly exhibits a sub-linear trend with $N$ and sigmoidal growth with respect to $N^{1-\alpha}$. Note also that the bottom left panel of Fig. 6(a) clearly shows that even at high task diversity values, relative distance slowly increases towards the memorizing predictor in a sigmoidal manner. This stands in contrast to Raventós et al. [25]’s claim that at high $D$ the Transformer only becomes closer to the generalizing predictor with increasing $N$ (see App. J for further discussion).
@@
  • Rapid behavior change near crossover. At the point where the two hypotheses have equal log-posterior odds, our model predicts there will be a crossover in which predictor dominates the posterior and explains model behavior. Given our sigmoidal functional form for $\eta(N,D)$, we can expect this change to be rapid—small variations in experimental conditions (e.g., task diversity) will yield large changes in model behavior. To confirm this, we plot the second derivative of the relative distance in Fig. 6(b), finding its magnitude is indeed greatest at the boundary $\eta(N,D)=0$.
@@
  • Scaling of time to Transience with Task-Diversity. For a given value of $D$, we can predict the critical training time at which transience, i.e., the crossover from generalizing to memorizing predictor, occurs by solving for $N^{*}$ in $\eta(N^{*},D)=0$ (i.e., relative distance $d_{\text{rel}}=0.5$). Specifically, we have: $N^{*}(D)=\left[\frac{\Delta K(D)^{\beta}}{\gamma\Delta L(D)}\right]^{\frac{1}{1-\alpha}}$. We thus plot the predicted time to transience as a function of $D$ in Fig. 6(c). As can be seen, beyond the two-hypotheses threshold (which determines the minimum $N$ value in which our model holds, see App. E), our predictions hold well with the empirically observed data. These results indicate super-linear growth of time to transience with task diversity. Importantly, we note that if this prediction derived from our model holds beyond our tested settings, it suggests that if the denominator in the expression becomes very small, e.g., when there are enough tasks in the mixture such that the memorizing and generalizing predictors produce similar outputs, the time to transience can approach infinity. In such a condition, generalization will persist regardless of how long a model is trained. Interestingly, we find that using a learning rate annealing schedule enhances the Transformer’s adherence to Bayes-optimal trajectories from generalization to memorization, and so we use it in this experiment. However, the effect of annealing appears to decay throughout training, yielding a slower rate of advancing toward memorization, which can be seen by the last two observed transience points veering from our model’s predictions (see App. K).
@@ 4.2 The Loss-Complexity Tradeoff
To further exemplify the intuition above, we can consider variations on the training pipeline that can be expected to affect a model’s ability to implement more complex predictors. Using the number of parameters as a control proxy to this end, we find that scaling MLP width raises the transition boundary between memorization and generalization—i.e., memorization is preferred more (Fig. 8). Our model closely captures this effect: we see a decreasing complexity penalty $\beta$ with larger MLP width, pointing toward the possibility that larger models penalize complexity less, yielding more memorization. Interestingly, we find that $\beta$ decays exponentially with MLP width. Finally, we also vary context length and task dimensionality, and find that, once again, our model captures the effects of these interventions (see App. H.4).
@@ 5 Discussion
Here, we discuss the relation between our work and other Bayesian or normative accounts of ICL. We provide an extended review of additional related works in App. B. Several prior works framed ICL as Bayesian at inference-time, meaning models implement a single Bayesian predictor in context [35, 73, 51, 74, 25, 62]. The focus on a single predictor in these works often led them to view ICL as a single strategy, with some focusing solely on a memorizing solution [35], while others refer mainly to a generalizing solution as ICL [25]. While there are cases in which in-context behavior is well approximated by a single predictor (e.g., the memorizing solution for low $N,D$), our results robustly show that several predictors are required to fully capture model behavior, and that the extent to which a particular predictor explains model behavior varies across training. Thus, a posterior-weighted average over different predictors, which considers a bias towards different predictors coming from training, rather than only inference-time, is required to fully capture model behavior across conditions. In contrast with studies focusing on inference-time, two recent works, Carroll et al. [30] and Elmoznino et al. [72], offer Bayesian or normative views of pretraining to perform ICL. Importantly, while these papers take different theoretical perspectives from ours, they arrive at a similar conclusion as us: the existence of a tradeoff between loss and solution complexity in pretraining. Elmoznino et al. [72] offer a normative theoretical analysis of training to perform ICL via next-token prediction loss, showing it yields an Occam’s razor objective which minimizes both loss and solution complexity. Carroll et al. [30] study task diversity effects and transient generalization in the linear regression setting of Raventós et al. [25]. Their Bayesian account of pretraining, which is rooted in theory of singular models [75] and makes different assumptions from ours, interestingly yields a relatively similar functional form for the posterior odds (though their form does not take into account neural scaling laws). However, their measure of complexity is architecture-dependent  [76], thus they only provide a qualitative analysis as they cannot directly estimate the complexity of Bayesian predictors. In contrast, our hierarchical Bayesian framework provides a quantitative, predictive account of pretraining phenomena, in addition to capturing inference-time behavior as a posterior-weighted average of solutions, which is not addressed by Carroll et al. [30] or Elmoznino et al. [72]. Despite that, we view these works as valuable contributions with complementary insights to ours, in particular regarding potential explanations for the source of neural networks’ simplicity bias.
@@
While we rely on a specific theoretical abstraction to arrive at our results, i.e., a hierarchical Bayesian model, we believe the predictive power of this abstraction corroborates its faithfulness. However, one limitation of our analysis is use of settings where model behavior is largely explained by only two predictors. It can be interesting to analyze settings where more predictors are feasible, e.g., the four predictors identified by Park et al. [26] in their task of learning a finite mixture of Markov chains. We believe this extension would be straightforward: we currently only characterize the dynamics of learning a memorizing vs. a generalizing predictor, but additionally modeling the statistics used to produce outputs (bigram vs. unigram) should accommodate the solutions identified by Park et al. [26]. Additionally, our current modeling analysis focuses on capturing Transformer behavior for sequences seen during training (ID), and extending it to OOD evaluation is an important step. Finally, a crucial limitation of our analysis comes from the simple relation we assume between algorithmic complexity and complexity of implementation by a Transformer—while the assumption of simplicity bias is well-backed by theoretical and empirical claims [46, 44, 43], we believe our assumed relation could be improved by building on recent advances defining architecture-dependent measures of how many effective parameters are used by a model to implement a solution [76, 77].
@@
We see several main takeaways from our work for understanding ICL, training dynamics, and generalization behavior in neural networks more broadly.
@@ Acknowledgments
We thank the Computation and Cognition Lab, in particular Ben Prystawski and Michael Li; Jay McClelland, Satchel Grant, Jerome Han and the PDP Lab; the Physics of Intelligence group at Harvard, especially Eric Bigelow and Sonia Murthy; the CRISP Lab at Harvard; Jesse Hoogland and Matthew Farrugia-Roberts; Surya Ganguli and the Neural Dynamics and Computation Lab; and Navin Goyal and the theory group at Microsoft Research India for useful discussions.
@@ Appendix A Glossary of Useful Terms
In-Context Learning (ICL). In this paper, we use a broad construal of ICL advanced by Lampinen et al. [37] and Olsson et al. [79]: any way in which models use their context to adapt their predictions and reduce loss can be considered as ICL. This notion of ICL encapsulates many forms of sequence modeling and instruction following, as well as the more traditional view of ICL as few-shot learning [2]. Hence, in this work, we use tasks that capture both sequence modeling [26] and few-shot learning [28, 25] notions of ICL. Generalizing Predictor ($G$). A predictor defined by a posterior-weighted average with a continuous prior over the true data-generating distribution $\mathcal{T}_{\text{true}}$. Such a predictor does not depend on the tasks seen during training, and can hence generalize well to novel, unseen tasks. This predictor maps onto ‘task-learning’ or ‘inference’ notions of ICL [26]. See App. G for the form of generalizing predictors in the studied settings. Memorizing Predictor ($M$). A predictor defined by a posterior-weighted average with a discrete prior defined over the distribution of tasks seen by the model during training, $\mathcal{T}_{train}$. Such a predictor depends on the tasks seen during training, and hence generalizes primarily to those tasks. This predictor maps onto ‘task-retrieval’ notions of ICL, though it also captures ‘in-weights learning’ in the classification task introduced by Chan et al. [20]. See App. G for the form of memorizing predictors in the studied settings. Task Diversity ($D$). For a mixture distribution $\mathcal{T}_{\text{train}}$ defined over $D$ tasks $\{f(\bm{w}_{1},\cdot),\dots,f(\bm{w}_{D},\cdot)\}$, $D$ is referred to as the task diversity of the mixture. Relative Distance ($d_{rel}$). A measure we use to characterize trade-off between predictors in settings where there are primarily two predictors (though one can easily generalize this measure to a setting with more predictors). Specifically, given a model $h$, two predictors $Q_{1},Q_{2}$, and a distance function $d$, we define the term $r:=\frac{d(h,Q_{1})-d(h,Q_{2})}{d(Q_{1},Q_{2})}$ and then relative distance as $d_{\text{rel}}=\nicefrac{{(r+1)}}{{2}}$. The latter operation rescales $r$ to a scale of 0 (if $h=Q_{1}$) to 1 (if $h=Q_{2}$), essentially assessing where the model $h$ lives when a line is drawn with endpoints ranging from $Q_{1}$ to $Q_{2}$. Posterior Odds, Prior Odds, Bayes Factor. Consider a set of observations $X$ and two hypotheses $H_{1}$ and $H_{2}$ that are being assessed as candidates to explain the data. Prior odds are defined as the ratio $\frac{P(H_{1})}{P(H_{2})}$, i.e., if no observations are seen yet, which hypothesis is apriori preferred. Bayes factor is defined as the ratio of likelihoods $\frac{P(X|H_{1})}{P(X|H_{2})}$, i.e., once the observations are received, we compare how likely individual hypotheses deem these observations. Finally, posterior odds are defined as the ratio $\frac{P(H_{1}|X)}{P(H_{2}|X)}=\frac{P(H_{1})}{P(H_{2})}\times\frac{P(X|H_{1})}{P(X|H_{2})}=\text{Prior odds}\times\text{Bayes factor}$, i.e., how do observations affect the prior to help assess which hypothesis is more likely. Rational Analysis. Rational analysis is a framework in cognitive science introduced by John R. Anderson [38], in which the behavior of a learner is explained as an optimal adaptation to its learning environment, given its goal and computational constraints. In the case of training a neural network, the learning environment is given by the training data distribution, the goal is given by the optimization objective, and the computational constraints can be seen as constraints and inductive biases of the architecture and optimization algorithm. The framework of rational analysis is characterized as normative, since it specifies how a learner should behave, in an optimal sense, given its environment, goals, and computational constraints. This approach has been successfully applied in cognitive science to explain a wide range of phenomena in humans [41, 39, 38, 80, 81]. Transient Generalization ($N^{*}$). For a given task diversity $D$, when a model is trained for sufficiently long, its behavior transitions from more closely resembling a generalizing predictor to more closely resembling a memorizing one. We define this point as the point at which the relative distance between memorizing and generalizing crosses $0.5$. This phenomenon is called transience, transient generalization, or the transient nature of in-context learning (see App. B for a longer discussion). The critical time to reach this point is denoted $N^{*}$ in the main paper.
@@ B.1 Prior Work Studying Task-Diversity Effects and Transience
We first discuss prior work studying the phenomenology of ICL that forms the core of our paper: task diversity effects and transience. These works often focus on providing mechanistic accounts that are more bottom-up in nature, i.e., suitable for studying specific settings. For example, the induction head is a core mechanism employed by a model to perform the in-context classification task by Singh et al. [24], Reddy [28]. When transience occurs, the role of induction head is diminished, since it is not vital for implementing a memorizing predictor. However, as we show, transience occurs across a spectrum of settings, including ones where induction heads are never learned during training (e.g., Balls and Urns, which we analyze using a one-layer Transformer)—prior mechanistic accounts would not help justify transience in such settings. Our work thus takes a top-down account, i.e., we offer insights based on a computational model of ICL developed by capturing its phenomenology. This helps identify relevant control variables controlling a dynamic we argue lies at the heart of transience and task diversity effects: tradeoff between loss and complexity of a solution, which manifests itself into transitions between generalizing and memorizing predictors. Reconciling our computational model with mechanistic approaches undertaken in prior work can be an exciting avenue for progress.
  • Task diversity Threshold. A phenomenon first popularized by Raventós et al. [25] in an in-context linear regression task, albeit originally demonstrated by Kirsch et al. [61] in a prior work on in-context classification of permuted MNIST images, and recently expanded to a Markov modeling setting by Park et al. [26], to another classification setting by Nguyen and Reddy [29], as well as to a modular arithmetic setting by He et al. [60]. Specifically, Raventós et al. [25] empirically show that if one trains Transformers on a mixture of linear regression tasks, there is a critical number of tasks below which the model behavior is well-characterized by a Bayesian predictor over the seen tasks (what we term the memorizing predictor), while above it the behavior is well-characterized by the standard solution of ridge regression. These results were expanded in a recent theoretical work by Lu et al. [82], who study the asymptotic dynamics of a linear attention Transformer and show the change in solution used by the model as a function of task diversity is a second-order phase transition. In contrast to their work, our empirical analysis covers a broad range of settings that involve sequence modeling, regression, and classification, while the analytical parts of our paper provide a predictive framework that identifies the relevant control variables and explains how they affect the behavior of standard, nonlinear Transformers across all studied settings.
@@
B.2 Hierarchical Bayesian Models of Learning to Learn
Hierarchical Bayesian models have been widely employed in cognitive science as models of ‘learning to-learn’, or meta-learning, in humans [83, 84, 85, 86, 87, 88]. These models allow the agent to learn the prior distribution from the data, rather than the prior being predefined. This allows learning inductive biases that constrain hypothesis spaces (called over-hypotheses), which can instruct future learning [87]. Additionally, it entails estimating evidence for several different hypothesis spaces (like the continuous vs. discrete priors in our work) and using these hypothesis spaces for generalization [86]. Note also that much of the work on hierarchical Bayesian models in cognitive science takes a normative perspective. Finally, we want to highlight the work of Grant et al. [89], who explicitly analyze a well-known meta-learning algorithm (MAML [90]) via a hierarchical Bayesian lens. We have been broadly influenced by this literature in approaching the study of ICL, which we believe has many direct parallels with work on learning to learn.
B.3 Broader Work on Understanding ICL